Accelerator Update for January 1, 2008 - January 7, 2008
Tuesday January 1
The midnight shift began with Operations monitoring the Tevatron (TeV) store 5819 and the Recycler stash, with the Antiproton Source stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking data. Operators tuned Booster and the Main Injector (MI) for increased efficiency and intensity.
During day shift operators reported that they had problems with the BNB autotune program. Operators contacted an expert who believed it was a MI8 autotune problem or due to bad Booster beam. He suggested we turn off BNB autotune program for about 20 minutes. Operators restored parameters for the program from a good file and adjusted the corrector magnet. Operations established store 5820 at 10:35 AM with an initial luminosity of 203.97E30. The Recycler stashed four times.
The evening shift operators reported that store 5280 collided all shift. The Recycler stashed three times and reached 300E10. There were problems with beam positions in Booster and MI; this was traced to the low level RF room air conditioning (AC) unit, which had its fan set to "ON" instead of "AUTO" and caused a temperature problem. Operators tuned and corrected some of the problems in Booster, but not in MI. Linac Klystron trips caused some downtime (the Booster beam position problems started after one of the Linac trips).
Wednesday January 2
The midnight shift began with Operations sending beam to NuMI and Sci/MiniBooNE. Operations established store 5821 at 5:09 AM with an initial luminosity of 181.6E30. The Recycler stashed twice during the shift.
The evening shift operators reported that store 5821 collided until it was terminated at 8 PM. D0 accessed the collision hall for 2.25 hours. During the access, the Recycler dampers went bad and caused the loss of about 200E10 from the stash. A further 10E10 was lost due to the Recycler tune program page that set all the quadrupole corrector power supplies to zero. A MI expert conducted some studies that halted running beam anywhere else. The Recycler stashed twice this shift. LRF4 tripped on a permanent inhibit; it took over an hour for it to recover.
Thursday January 3
The midnight shift began with the D0 experimenters in access. The Recycler had problems with transverse emittances, which were too big to transfer antiprotons from the Accumulator. Operators were trying to slowly squeeze the Recycler beam to make room. The stash was 117E10 and the stack was 85E10. Operations sent beam to NUMI and Sci/MiniBooNE when it was available. Per a Linac expert request, operators adjusted the filament for the LRF4 power amplifier (PA) to make it stable. Operations established store 5824 at 6:55 AM with an initial luminosity of 155.6E30. The Recycler stashed twice during this shift.
The day shift operators reported that store 5824 collided all shift. The Recycler stashed twice during this shift. An Accumulator kicker (A:IKIK) failed requiring an access to replace a matching capacitor. Linac experts replaced a PA for a Linac RF station (LRF4). Operators switched to the H- Source due to vacuum problems with the I- Source. Efforts to pump on the PA vacuum were underway by the end of shift. MI experts completed a study just before the beam was turned off.
The evening shift operators reported that store 5824 collided all shift. Linac resumed sending beam at 4:30 PM, which was after the switch to the H- Source. Linac experts tuned for an hour and then Booster experts tuned. Operators reported that the TeV sector A4 vacuum was degrading. The Recycler stashed four times during the shift.
Friday January 4
The midnight shift operators reported that store 5824 collided all shift. Operations sent beam to NuMI and Sci/MiniBooNE. TeV experts began a one-hour end-of-store study at 6:30 AM; Operations terminated the store at 7:30 AM. The Recycler stashed four times. Operators tuned Booster for increased intensity.
The day shift operators reported they had established store 5825 at 9:49 AM with an initial luminosity of 232.74E30. Booster experts continued to tune up for increased intensities and efficiencies. Recycler experts conducted diffusion studies. The Recycler stashed twice during this shift. On the second stash there was a problem that required a return to stacking to investigate a transfer problem. MI experts conducted a collimator study and a 11 batch slipstacking study.
The evening shift operators reported that store 5825 collided all shift. Operators reported that they had restarted the BNB autotune program. The Recycler stashed five times.
Saturday January 5
The midnight shift operators reported that store 5825 collided until it was terminated at 6:46 AM. D0 made an access to their collision hall. A Booster kicker (MKS07) started sparking; experts repaired a PFN cable. The Recycler stashed three times. Operations sent beam to NuMI and Sci/MiniBooNE.
The day and evening shifts began with D0 coming out of access. Operations put the TeV through a dry squeeze and then established store 5827 at 1:28 PM with an initial luminosity of 231.38E30. The Recycler stashed four times. Operators reported the Pbar orbit changed, which they corrected by adjusting a Lambertson (I:LAM52).
Sunday January 6
The midnight shift operators reported that store 5827 collided until 7:36 AM, when it was lost due to an antiquench of C:DQ0; it was a clean abort with no quench. Operators investigated and then consulted experts. The Recycler stashed six times. Various Linac trips caused downtime. Operations sent beam to NuMI and Sci/MiniBooNE.
The day and evening shift operators reported at 10 AM that experts had repaired C:DQ0. During shot setup a power supply’s (T:VD49) fuse faulted, it was replaced and the TeV shot setup started again at 1 PM. Operations established store 5830 at 2:45 PM with an initial luminosity of 232.88E30. There were problems with this store: the T:BEAM readbacks caused about 45 min of downtime and operators informed experts; the TeV antiproton injection closure did not work; and a collimator (T:D173DH) did not go all the way in and required a rescrape. The Recycler stashed four times. The Accumulator extraction shutter (A:ESHUT) controller failed. Operators investigated, informed experts, and determined that the shutter would remain in the open position until Monday morning.
Monday January 7
The midnight shift operators reported that store 5830 collided all shift. The Recycler stashed six times during this shift. Operators tuned the Linac and Booster to correct beam position problems with only a small amount of success. A TeV separator (C49V) sparked and affected the luminosity. Operators reported a MI RF watchdog trip due to vacuum problems in the RF region. Operations sent beam to NuMI and Sci/MiniBooNE when it was available.
The Shift Summary for January 7 to 9 will be on Thursday
Comments and Suggestions
|last modified 01/08/2008 email Fermilab|
|Security, Privacy, Legal|